OpenStudy (anonymous):

solve by factoring: 4x^2-9=0

OpenStudy (anonymous):

\[(2x)^2 - 3^2 = 0\] Can you use the formula now???

OpenStudy (anonymous):

what formula would i use?

OpenStudy (anonymous):

a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

\[(a^2 - b^2) = (a-b)(a+b)\]

OpenStudy (anonymous):

is it (2x-3)(2x+3)=0

OpenStudy (anonymous):

Yes dear, you are right.. Now put separately both the factors = 0 (2x - 3) = 0 and (2x + 3) = 0 and now find x from here..